I have the equations x + y = 56 and x = 3y = 56
x is one floorplan of a house and y is the other floorplan
There are 3 times as many houses with the floorplan y.
Need to solve for x and y.
I have tried to substitute without success!
I come up with x = 15 and y would equal 15 * 3 = 45 10 + 45 = 55 not 56!

Answer by jorel1380(3719)

If x + y = 56, and x=3y, then
x+y=56
3y+y=56
4y=56
y=14
x=3y=42
There are 14 houses with floorplan y. There are 42 houses with floorplan x.